Compute the sample variance and standard deviation for the returns: R1 = 15%, R2 = 20.65%, R3 = -13.51%, R4 = -16.67%
Answer #1

Expected Return = [=1R(i)] / n

= [15% + 20.65% + (-13.51%) + (-16.67%)] / 4 = 5.47% / 4 = 1.3675%

Sample Variance = [=1{Expected Return - R(i)}2] / (n - 1)

= [(1.3675% - 15%)2 + (1.3675% - 20.65%)2 + (1.3675% + 13.51%)2 + (1.3675% + 16.67%)2] / (4 - 1)

= [185.85%2 + 371.81%2 + 221.34%2 + 325.35%2] / 3 = 1,104.35%2 / 3 = 368.12%2

Sample Standard Deviation = [Sample Variance]1/2

= [368.12%2]1/2 = 19.19%